OpenStudy (anonymous):

I have a GRE question: How many different 3-digit numbers can be formed, using the digits from 0 to 9 inclusive, if no digit is used more than once? A 504 B 648 C 729 D 900 E 1,000

OpenStudy (anonymous):

Do you think it would be easier to count the number of 3 digit numbers with repeats, or the number of 3 digit numbers with no repeats?

OpenStudy (anonymous):

This is actually just a simple permutation task, isn't it?

OpenStudy (anonymous):

Including repeats the answer is 1,000 but the question says "no repeats" Something like 10!/7! is what I figure the answer should be, but given the 5 solutions, only option C comes close.

OpenStudy (anonymous):

I see, you can't use 0 as a leading digit, that is what makes it complicated. I would start with 10 P 7, then try to remove the leading 0 numbers.

OpenStudy (anonymous):

We know with 10 P 7, we won't have anything like 001, But we will have something like 012

OpenStudy (anonymous):

Okay I've been saying 10 P 7, I mean 10 P 3

OpenStudy (anonymous):

Assuming the first digit is 0 though, the remaining qualifiers are 9 P 2

OpenStudy (anonymous):

So then you have 10 P 3 - 9 P 2 = 10!/7! - 9!/7! =10 * 9!/7! - 9!/7! = 9*9!/7!

OpenStudy (anonymous):

@Matt.Mawson I end up getting an answer that is listed.